# Learning Equilibria from Data: Provably Efficient Multi-Agent Imitation Learning

### v1

@@ Abstract
This paper provides the first expert sample complexity characterization for learning a Nash equilibrium from expert data in Markov Games. We show that a new quantity named the single policy deviation concentrability coefficient is unavoidable in the non-interactive imitation learning setting, and we provide an upper bound for behavioral cloning (BC) featuring such coefficient. BC exhibits substantial regret in games with high concentrability coefficient, leading us to utilize expert queries to develop and introduce two novel solution algorithms: MAIL-BRO and MURMAIL. The former employs a best response oracle and learns an $\varepsilon$-Nash equilibrium with $\mathcal{O}(\varepsilon^{-4})$ expert and oracle queries. The latter bypasses completely the best response oracle at the cost of a worse expert query complexity of order $\mathcal{O}(\varepsilon^{-8})$. Finally, we provide numerical evidence, confirming our theoretical findings.
Learning Equilibria from Data:
Provably Efficient Multi-Agent Imitation Learning
Till Freihaut*
freihaut@ifi.uzh.ch
University of Zurich
Luca Viano*
luca.viano@epfl.ch
EPFL
Volkan Cevher
volkan.cevher@epfl.ch
EPFL
Matthieu Geist
matthieu@earthspecies.org
Earth Species Project
Giorgia Ramponi
ramponi@ifi.uzh.ch
University of Zurich
$\ast$\$$\ast$\$footnotetext: Equal contribution.
@@ 1 Introduction
In this work, we consider 2-player Zero-Sum Markov Games [^1] For the sake of simplicity, the main text will focus on this case. The appendix outlines the extension to $n$ players’ general sum games. where the agents’ rewards are perfectly opposing, i.e., $r_{1}(s,a,b)=-r_{2}(s,a,b)$. In this setting, denoting the state value function of a strategy pair $\mu,\nu$ as $V^{\mu,\nu}:{\mathcal{S}}\rightarrow\mathbb{R}$, we measure the gap of a strategy pair to an NE by the following metric
@@
where $s_{0}$ is the starting state of the game [^2] We will relax this to a stochastic starting state in the next sections. and $\nu^{\star}$ denotes one of the strategies from the set of best responding strategies to $\mu$. That is, $\mu^{\star}\in\mathrm{br}(\nu):=\mathop{\rm argmax}_{\mu}V^{\mu,\nu}(s_{0})$ and $\nu^{\star}\in\mathrm{br}(\mu):=\mathop{\rm argmin}_{\nu}V^{\mu,\nu}(s_{0}).$ This objective has been widely adopted in Multi-Agent Reinforcement Learning (see, e.g., (Cui and Du, 2022a, b)) and it is easily motivated by the fact that any strategy profile output by an algorithm under study $(\mu_{\mathrm{out}},\nu_{\mathrm{out}})$ such that $\mathrm{Nash}\text{-}\mathrm{Gap}(\mu_{\mathrm{out}},\nu_{\mathrm{out}})\leq\varepsilon$ is an $\epsilon$-approximate Nash equilibrium, often shortened as $\varepsilon$-NE. However, it remained largely unexplored in the MAIL setting until the seminal work of Tang et al. (2024), who showed that minimizing the Nash Gap is fundamentally hard in MAIL since deviations in out-of-distribution states can incur linear regret.
Table 1: For simplicity, we report results for the two players zero sum with discount factor $\gamma$, finite state space $\left|{{\mathcal{S}}}\right|$, finite action spaces ${\mathcal{A}}$, $\mathcal{B}$. Let $\left|{{\mathcal{A}}_{\mathrm{max}}}\right|=\max{\left|{{\mathcal{A}}}\right|,\left|{\mathcal{B}}\right|}$. Being consistent with Tang et al. (2024), we denote $\beta=\min_{s\in{\mathcal{S}}}d^{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}(s)$, by $u$ the recoverability coefficient and with $H$ the finite horizon of the considered game. Moreover, we refer to Tang et al. (2024) for the definition of the convex functions $\ell_{\mathrm{MALICE}}$ and $\ell_{\mathrm{BLADES}}$. Additionally, for the behavioral cloning (BC) output pair $\widehat{\mu},\widehat{\nu}$ with an input dataset $\mathcal{D}$ we define $\mathcal{C}(\hat{\mu},\hat{\nu})=\max_{\mu\in\mathrm{br}({\widehat{\nu}})}\left\|\frac{d^{\mu,\nu^{\operatorname{E}}}}{d^{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}}\right\|_{\infty}+\max_{\nu\in\mathrm{br}({\widehat{\mu}})}\left\|\frac{d^{\mu^{\operatorname{E}},\nu}}{d^{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}}\right\|_{\infty}$. For this comparison, notice that the main text by Tang et al. (2024) focuses on learning correlated equilibria, but as specified in their appendix, the same proofs can be performed for the problem of learning Nash equilibria. For the algorithms presented by Tang et al. (2024), we can not specify the bound on the number of expert queries since their analysis as an error propagation only flavor. As a final minor difference, we apply our analysis to the infinite horizon discounted setting, which is more relevant for practical settings. Finally, we abbreviated Queriable Expert by QE.
@@
BC Tang et al. (2024) $\beta>0$ 00 ( analytical solution is available) $\mathcal{O}\left({uH\varepsilon\beta^{-1}}\right)$ Not specified $\epsilon$-accurate TV minimizer
@@
BC (Our analysis) $\mathcal{C}(\hat{\mu},\hat{\nu})<\infty$ 00 ( analytical solution is available) $\mathcal{O}\left({\varepsilon}\right)$ $\widetilde{\mathcal{O}}\left({\frac{\left|{{\mathcal{S}}}\right|\left|{{\mathcal{A}}_{\mathrm{max}}}\right|\mathcal{C}(\hat{\mu},\hat{\nu})}{(1-\gamma)^{4}\varepsilon^{2}}}\right)$ None
@@
    We provide a sample complexity analysis for BC, revealing the emergence of a single deviation concentrability coefficient (Theorem 3.1).
@@
    We formally separate MAIL from SAIL, proving in Theorem 3.2 that even with fully known transitions, for any non-interactive imitation learning algorithm (like BC) there exists a Markov Game with infinite single deviation concentrability coefficient where the Nash Gap remains constant even with infinite expert data.
@@
    On the positive side, we show that the dependence on the concentrability coefficient can be avoided if an interactive expert is available. In particular, assuming access to a Best Response Oracle, we propose an algorithm that achieves an $\epsilon$-NE with $\mathcal{O}(\epsilon^{-4})$ expert queries and oracle calls (Algorithm 2).
@@
    Additionally, we develop an algorithm that avoids the Best Response oracle and the concentrability coefficient simultaneously, achieving an $\epsilon$-NE with $\mathcal{O}(\epsilon^{-8})$ expert queries. Moreover, the algorithm is computationally efficient. Its design is based on the novel principle of maximum uncertainty response.
For clarity, we report a comparison of our results with existing MAIL algorithms in Table 1.
@@ 2 Preliminaries
Additionally, we define the state visitation probability induced by a policy pair $(\mu,\nu)$ as $d^{\mu,\nu}(s^{\prime}):=(1-\gamma)\mathbb{E}_{\mu,\nu}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbf{1}_{\{S_{t}=s^{\prime}\}}\mid s_{0}\sim d_{0}\right]$ If one player’s policy is fixed, then the Markov Game induces a Markov decision process (MDP). Assuming that player 2 fixes their strategy, the induced transition function for a given state-action pair $(s,a)\in{\mathcal{S}}\times{\mathcal{A}}$ to new state $s^{\prime}\in{\mathcal{S}}$ is given by $P_{\nu}(s^{\prime}\mid s,a):=\sum_{b\in{\mathcal{B}}}\nu(b\mid s)P(s^{\prime}\mid s,a,b).$ It is analogously defined if the policy of player 1 is fixed. Additionally, for a fixed strategy of the opponent player, we define the best response set as $\mathrm{br}(\nu)=\mathop{\rm argmax}_{\mu\in\Pi}\left\langle{d_{0}},{V^{\mu,\nu}}\right\rangle$ and $\mathrm{br}(\mu)=\mathop{\rm argmin}_{\nu\in\Pi}\left\langle{d_{0}},{V^{\mu,\nu}}\right\rangle,$ respectively, where $\Pi$ denotes the set of all possible policies and $\mu^{\star},\nu^{\star}$ as elements of these sets. It is important to note that the best response may not be unique, but the value is. A pair of policies is called a Nash equilibrium if both policies are best responses to each other. Last, we introduce the Nash gap, which measures how close a given policy pair $(\mu,\nu)$ is to a NE:
$\mathrm{Nash}\text{-}\mathrm{Gap}(\mu,\nu):=\left\langle{d_{0}},{V^{\mu^{\star},\nu}-V^{\mu,\nu^{\star}}}\right\rangle.$ (1)
The Nash-Gap has the desirable property, that $\mathrm{Nash}\text{-}\mathrm{Gap}(\mu,\nu)=0,$ if $(\mu,\nu)$ is a NE and otherwise $\mathrm{Nash}\text{-}\mathrm{Gap}(\mu,\nu)>0$.
@@
In non-interactive MAIL, the learner observes a dataset ${\mathcal{D}}:=\{\tau_{k}\}_{k=1}^{N}$ containing $N$ trajectories collected in the two-player zero-sum Markov Game, where the actions are sampled from the NE expert policy pair $(\mu^{\operatorname{E}},\nu^{\operatorname{E}}).$ For each trajectory $\tau_{k}$, a random length $H\sim\mathrm{Geo}(1-\gamma)$ is sampled and then the sequence of states and (joint-)actions up to time $H$ are saved, i.e. $\tau_{k}:=\{(s_{t},a_{t},b_{t})\}_{t=1}^{H}.$ After such dataset is collected, the learner can no longer collect new expert data. For this reason, we refer to the setting as non-interactive. Moreover, the learner might know the transition function of the Markov Game. The learner’s goal is to adopt an algorithm $\mathrm{Alg}$ that takes as input $\mathcal{D}$, and outputs a pair of policies $(\widehat{\mu},\widehat{\nu})$ such that $\mathbb{E}_{\mathrm{Alg}}\left[{\mathrm{Nash}\text{-}\mathrm{Gap}(\widehat{\mu},\widehat{\nu})}\right]<\varepsilon.$
@@
In interactive MAIL, there is no initial dataset $\mathcal{D}$. The learner interacts with the environment for a certain number of rounds. At each round, the learner can collect a trajectory with a chosen policy pair and decide to query the expert at the visited states. The learner’s goal is to adopt an algorithm $\mathrm{Alg}$ that after $\mathrm{poly}(\varepsilon^{-1})$ main expert queries outputs a pair of policies $(\widehat{\mu},\widehat{\nu})$ such that $\mathbb{E}_{\mathrm{Alg}}\left[{\mathrm{Nash}\text{-}\mathrm{Gap}(\widehat{\mu},\widehat{\nu})}\right]<\varepsilon.$ Compared to the non-interactive setting, the expert can be queried during learning.
@@ 3 On the sample complexity of Multi-Agent Behavior Cloning
Interestingly, our upper bound depends on a novel quantity $\mathcal{C}:\Pi\times\Pi\rightarrow\mathbb{R}$ dubbed single policy deviation concentrability coefficient, which is an infinite norm ratio between the occupancy distributions related to the notion of data coverage assumptions needed in Offline Zero-Sum Markov Games (Cui and Du, 2022a; Zhong et al., 2022) and concentrability coefficients in approximate dynamic programming (Scherrer et al., 2012; Geist et al., 2019; Vieillard et al., 2020). Contrary to the analysis of Tang et al. (2024), we do not require that the occupancy measure of the equilibrium policy pair used to collect $\mathcal{D}$ is lower bounded by $\beta$. Therefore, our analysis also applies to the realistic setting where some states have zero probability to appear in $\mathcal{D}$.
@@
Behavioural cloning in Markov Games.
In the context of Markov games, BC aims to recover a pair of policies $(\widehat{\mu},\widehat{\nu})$ from expert demonstrations ${\mathcal{D}}$ based on maximum likelihood estimation. Formally, we have that $(\widehat{\mu},\widehat{\nu})=\mathop{\rm argmax}_{(\mu,\nu)}\sum_{\tau\in{\mathcal{D}}}\log\left({\mathbb{P}}(\tau;\mu,\nu)\right),$ where ${\mathbb{P}}(\tau;\mu,\nu)=d_{0}(s_{0})\prod_{h=1}^{H}\mu(a\mid s)\nu(b\mid s)P(s^{\prime}\mid s,a,b)$ is the probability of generating trajectory $\tau$ under policies $(\mu,\nu)$, where $H\sim\mathrm{Geo}(1-\gamma)$. In the tabular set-up, we can obtain the closed-form solution of the above optimization problem $\widehat{\mu}(a\mid s)=\frac{N(s,a)}{N(s)},$ if $N(s)>0$ and $\widehat{\mu}(a\mid s)=\frac{1}{\left|{{\mathcal{A}}}\right|}$ otherwise. Similarly, this holds for $\widehat{\nu}(b\mid s)$ by replacing $N(s,a)$ with $N(s,b)$. Here $N(s,a),N(s,b)$ and $N(s)$ denote the number of times that state-action pair $(s,a),(s,b)$ and state $s$ appear in ${\mathcal{D}}$.
Now, we can state our result for the upper bound of Behavior Cloning when minimizing the Nash Gap (1). We give a proof sketch below the theorem and the full proof can be found in Appendix D.
@@
Let $(\mu^{E},\nu^{E})$ denote a Nash equilibrium policy pair in a two-player zero-sum Markov game, and let ${\mathcal{D}}$ contain trajectories from this expert policy pair. Let $(\widehat{\mu},\widehat{\nu})$ be the policies obtained via Behavior Cloning from ${\mathcal{D}}$ of size $N$. Then, with probability at least $1-\delta$, it holds:
$\mathrm{Nash}\text{-}\mathrm{Gap}(\widehat{\mu},\widehat{\nu})\leq{\mathcal{C}}(\widehat{\mu},\widehat{\nu})\frac{8}{(1-\gamma)^{2}}\sqrt{\frac{\left|{{\mathcal{S}}}\right|\left|{{\mathcal{A}}_{\max}}\right|\log^{2}(2\left|{{\mathcal{S}}}\right|/\delta)}{N}},$
where $\mathcal{C}(\hat{\mu},\hat{\nu}):=\max_{\mu\in\mathrm{br}({\widehat{\nu}})}\left\|\frac{d^{\mu,\nu^{\operatorname{E}}}}{d^{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}}\right\|_{\infty}+\max_{\nu\in\mathrm{br}({\widehat{\mu}})}\left\|\frac{d^{\mu^{\operatorname{E}},\nu}}{d^{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}}\right\|_{\infty}.$
@@
where $\mu^{\star}\in\mathrm{br}(\widehat{\nu}),\nu^{\star}\in\mathrm{br}(\widehat{\mu}).$ Next, we can upper bound the two error terms separately. Note that the error terms each share one fixed policy, therefore, we can apply a version of the performance difference lemma, the triangle inequality, and fix one best response for each player to obtain
@@
Last, we do a change of measure to get the expectation with respect to the expert policy pair. Then, we bound the ratio of the state visitation distribution and bound the expectation with concentration inequalities to obtain with probability of at least $1-\delta$
$V^{\mu^{\star},\widehat{\nu}}(s_{0})-V^{\mu^{\star},\nu^{\operatorname{E}}}(s_{0})\leq\frac{8}{(1-\gamma)^{2}}\max_{\mu\in\mathrm{br}(\widehat{\nu})}\left\|\frac{d^{\mu,\nu^{\operatorname{E}}}}{d^{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}}\right\|_{\infty}\sqrt{\frac{\left|{{\mathcal{S}}}\right|\left|{{\mathcal{B}}}\right|\log^{2}(2\left|{{\mathcal{S}}}\right|/\delta)}{N}}.$
Analogous calculations for the second player complete the proof. The full proof is given in Appendix D. ∎
We now discuss several important implications of the derived theorem, particularly focusing on the quantity $\mathcal{C}(\hat{\mu},\hat{\nu})$, referred to as the single policy deviation concentrability coefficient (see, e.g., (Cui and Du, 2022a; Zhong et al., 2022)). Intuitively, the theorem indicates that if the best response of the recovered policy shifts the support of the state visitation distribution away from the one induced by the observed Nash equilibrium, the corresponding objective becomes unbounded.
@@
While restricting, this requirement is weaker than a uniform lower bound on the equilibrium state occupancy measure assumed by Tang et al. (2024), that is $d^{\mu_{E},\nu_{E}}\geq\beta$. In particular, it always holds that $\mathcal{C}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})\leq\beta^{-1}$.
On the positive side, we can notice that $\mathcal{C}(\hat{\mu},\hat{\nu})$ equals $\mathcal{C}(\mu^{\operatorname{E}},\nu^{\operatorname{E}}):=\max_{\mu\in\mathrm{br}(\nu^{\operatorname{E}})}\left\|{\frac{d^{\mu,\nu^{\operatorname{E}}}}{d^{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}}}\right\|_{\infty}+\max_{\nu\in\mathrm{br}(\mu^{\operatorname{E}})}\left\|{\frac{d^{\mu^{\operatorname{E}},\nu}}{d^{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}}}\right\|_{\infty}$ in the limit of infinite data in the dataset $\mathcal{D}$. Moreover, $\mathcal{C}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})=1$ if $\mu^{\operatorname{E}}$ is the unique best response to $\nu^{\operatorname{E}}$ and vice versa. It follows that BC is expected to work well under this condition which, for example, can be achieved in entropy regularized games.
On the negative side, we show that there exists a zero-sum Markov game in which $\mathcal{C}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$ is unbounded, and we show that in such a game, no non-interactive algorithm can recover a Nash profile even under an infinite amount of data. We present this result in the next section.
@@
3.1 Necessity of $\mathcal{C}(\mu,\nu)$ in non-interactive MAIL
In this section, we provide the negative result, that a Markov Game exists, such that the single deviation concentrability coefficient of Theorem 3.1 is unbounded.
The first hardness results to minimize the Nash Gap in Multi-Agent Imitation Learning were derived by  Tang et al. (2024, Thm. 4.3). Next, we will give a stronger result, showing that even in the case of full knowledge of the transition model and perfect recovery of the state visitation distribution of the expert, the Nash gap is of the order $(1-\gamma)^{-1}$. A detailed discussion on the difference between the following result and the one obtained by Tang et al. (2024, Thm. 4.3) can be found in Appendix J. An illustration of the Zero-Sum Markov game can be found in Fig. 1 and the full proof in Appendix E.

### v2

@@ Abstract
This paper provides the first expert sample complexity characterization for learning a Nash equilibrium from expert data in Markov Games. We show that a new quantity named the all policy deviation concentrability coefficient is unavoidable in the non-interactive imitation learning setting, and we provide an upper bound for behavioral cloning (BC) featuring such coefficient. BC exhibits substantial regret in games with high concentrability coefficient, leading us to utilize expert queries to develop and introduce two novel solution algorithms: MAIL-BRO and MURMAIL. The former employs a best response oracle and learns an $\varepsilon$-Nash equilibrium with $\mathcal{O}(\varepsilon^{-4})$ expert and oracle queries. The latter bypasses completely the best response oracle at the cost of a worse expert query complexity of order $\mathcal{O}(\varepsilon^{-8})$. Finally, we provide numerical evidence, confirming our theoretical findings.
[…]
@@ 1 Introduction
In this work, we consider 2-player Zero-Sum Markov Games [^1] For the sake of simplicity, the main text will focus on this case. The appendix outlines the extension to $n$ player general-sum games. where the agents’ rewards are perfectly opposing, i.e., $r_{1}(s,a,b)=-r_{2}(s,a,b)$. In this setting, denoting the state value function of a strategy pair $\mu,\nu$ as $V^{\mu,\nu}:{\mathcal{S}}\rightarrow\mathbb{R}$, we measure the gap of a strategy pair to an NE by the following metric
@@
where $s_{0}$ is the starting state of the game [^2] We will relax this to a stochastic starting state in the next sections. and $\nu^{\star}$ denotes one of the strategies from the set of best-response strategies to $\mu$. That is, $\mu^{\star}\in\mathrm{br}(\nu):=\mathop{\rm argmax}_{\mu}V^{\mu,\nu}(s_{0})$ and $\nu^{\star}\in\mathrm{br}(\mu):=\mathop{\rm argmin}_{\nu}V^{\mu,\nu}(s_{0}).$ This objective has been widely adopted in Multi-Agent Reinforcement Learning (see, e.g., (Cui and Du, 2022a, b)) and it is easily motivated by the fact that any strategy profile output by an algorithm under study $(\mu_{\mathrm{out}},\nu_{\mathrm{out}})$ such that $\mathrm{Nash}\text{-}\mathrm{Gap}(\mu_{\mathrm{out}},\nu_{\mathrm{out}})\leq\varepsilon$ is an $\varepsilon$-approximate Nash equilibrium, often shortened as $\varepsilon$-NE. However, it remained largely unexplored in the MAIL setting until the seminal work of Tang et al. (2024), who showed that minimizing the Nash Gap is fundamentally hard in MAIL since deviations in out-of-distribution states can incur linear regret.
Table 1: For simplicity, we report results for the two-player zero-sum with discount factor $\gamma$, finite state space $\left|{{\mathcal{S}}}\right|$, finite action spaces $\mathcal{A}$}, $\mathcal{B}$. Let $\left|{{\mathcal{A}}_{\mathrm{max}}}\right|=\max{\left|{{\mathcal{A}}}\right|,\left|{\mathcal{B}}\right|}$. Being consistent with Tang et al. (2024), we denote $\beta=\min_{s\in{\mathcal{S}}}d^{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}(s)$, by $u$ the recoverability coefficient and with $H$ the finite horizon of the considered game. Moreover, we refer to Tang et al. (2024) for the definition of the convex functions $\ell_{\mathrm{MALICE}}$ and $\ell_{\mathrm{BLADES}}$. Additionally, for the behavioral cloning (BC) output pair $\widehat{\mu},\widehat{\nu}$ with an input dataset $\mathcal{D}$ we define $\mathcal{C}_{\max}=\max_{\mu,\nu}\max\left\{\max_{\nu^{\star}\in\mathrm{br}(\mu)}\left\|{\frac{d^{\mu^{\operatorname{E}},\nu^{\star}}}{\rho}}\right\|_{\infty},\max_{\mu^{\star}\in\mathrm{br}(\nu)}\left\|{\frac{d^{\mu^{\star},\nu^{\operatorname{E}}}}{\rho}}\right\|_{\infty}\right\},$ where $\rho$ is any state state distribution, in BC we set it to $\rho=d^{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}.$ For this comparison, notice that the main text by Tang et al. (2024) focuses on learning correlated equilibria, but as specified in their appendix, the same proofs can be performed for the problem of learning Nash equilibria. For the algorithms presented by Tang et al. (2024), we can not specify the bound on the number of expert queries since their analysis as an error propagation only flavor. As a final minor difference, we apply our analysis to the infinite horizon discounted setting, which is more relevant for practical settings. Finally, we abbreviated Queriable Expert by QE.
@@
BC Tang et al. (2024) $\beta>0$ 0 (analytical solution is available) $\mathcal{O}\left({uH\varepsilon\beta^{-1}}\right)$ Not specified $\varepsilon$-accurate TV minimizer
@@
BC (Our analysis) $\mathcal{C}_{\max}<\infty$ 0 (analytical solution is available) $\mathcal{O}\left({\varepsilon}\right)$ $\widetilde{\mathcal{O}}\left({\frac{\left|{{\mathcal{S}}}\right|\left|{{\mathcal{A}}_{\mathrm{max}}}\right|\mathcal{C}^{2}_{\max}}{(1-\gamma)^{4}\varepsilon^{2}}}\right)$ None
@@
    We provide a sample complexity analysis for BC, revealing the emergence of an all deviation concentrability coefficient (Theorem˜3.1).
@@
    We formally separate MAIL from SAIL, proving in Theorem˜3.2 that even with fully known transitions, for any non-interactive imitation learning algorithm (like BC) there exists a Markov Game with infinite single deviation concentrability coefficient where the Nash Gap remains constant even with infinite expert data.
@@
    On the positive side, we show that the dependence on the concentrability coefficient can be avoided if an interactive expert is available. In particular, assuming access to a Best Response Oracle, we propose an algorithm that achieves an $\varepsilon$-NE with $\mathcal{O}(\varepsilon^{-4})$ expert queries and oracle calls (Algorithm˜2).
@@
    Additionally, we develop an algorithm that avoids the Best Response oracle and the concentrability coefficient simultaneously, achieving an $\varepsilon$-NE with $\mathcal{O}(\varepsilon^{-8})$ expert queries. Moreover, the algorithm is computationally efficient. Its design is based on the novel principle of maximum uncertainty response.
For clarity, we report a comparison of our results with existing MAIL algorithms in Table˜1.
@@ 2 Preliminaries
Additionally, we define the state visitation probability induced by a policy pair $(\mu,\nu)$ as $d^{\mu,\nu}(s^{\prime}):=(1-\gamma)\mathbb{E}_{\mu,\nu}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbf{1}_{\{S_{t}=s^{\prime}\}}\mid s_{0}\sim d_{0}\right]$ . If one player’s policy is fixed, then the Markov Game induces a Markov decision process (MDP). Assuming that player 2 fixes their strategy, the induced transition function for a given state-action pair $(s,a)\in{\mathcal{S}}\times{\mathcal{A}}$ to new state $s^{\prime}\in{\mathcal{S}}$ is given by $P_{\nu}(s^{\prime}\mid s,a):=\sum_{b\in{\mathcal{B}}}\nu(b\mid s)P(s^{\prime}\mid s,a,b).$ It is analogously defined if the policy of player 1 is fixed. Additionally, for a fixed strategy of the opponent player, we define the best response set as $\mathrm{br}(\nu)=\mathop{\rm argmax}_{\mu\in\Pi}\left\langle{d_{0}},{V^{\mu,\nu}}\right\rangle$ and $\mathrm{br}(\mu)=\mathop{\rm argmin}_{\nu\in\Pi}\left\langle{d_{0}},{V^{\mu,\nu}}\right\rangle,$ respectively, where $\Pi$ denotes the set of all possible policies and $\mu^{\star},\nu^{\star}$ as elements of these sets. It is important to note that the best response may not be unique, but the value is. A pair of policies is called a Nash equilibrium if both policies are best responses to each other. Last, we introduce the Nash gap, which measures how close a given policy pair $(\mu,\nu)$ is to a NE:
$\mathrm{Nash-}\mathrm{Gap}(\mu,\nu):=\left\langle{d_{0}},{V^{\mu^{\star},\nu}-V^{\mu,\nu^{\star}}}\right\rangle.$ (1)
The Nash-Gap has the desirable property, that $\mathrm{Nash-}\mathrm{Gap}(\mu,\nu)=0,$ if $(\mu,\nu)$ is a NE and $\mathrm{Nash-}\mathrm{Gap}(\mu,\nu)>0,$ otherwise.
@@
[…]
@@
[…]
@@ 3 On the sample complexity of Multi-Agent Behavior Cloning
Interestingly, our upper bound depends on a novel quantity $\mathcal{C}_{\max}\in\mathbb{R}$ dubbed all policy deviation concentrability coefficient, which is an infinite norm ratio between the occupancy distributions related to the notion of data coverage assumptions needed in Offline Zero-Sum Markov Games (Cui and Du, 2022a; Zhong et al., 2022) and concentrability coefficients in approximate dynamic programming (Scherrer et al., 2012; Geist et al., 2019; Vieillard et al., 2020). Contrary to the analysis of Tang et al. (2024), we do not require that the occupancy measure of the equilibrium policy pair used to collect $\mathcal{D}$ is lower bounded by $\beta$. Therefore, our analysis also applies to the realistic setting where some states have zero probability to appear in $\mathcal{D}$.
@@
Behavioral cloning in Markov Games.
[…]
Now, we can state our result for the upper bound of Behavior Cloning when minimizing the Nash Gap (1). We give a proof sketch below the theorem and the full proof can be found in Appendix˜C.
@@
[…]
$\mathrm{Nash-}\mathrm{Gap}(\widehat{\mu},\widehat{\nu})\leq{\mathcal{C}}_{\max}\frac{8}{(1-\gamma)^{2}}\sqrt{\frac{\left|{{\mathcal{S}}}\right|\left|{{\mathcal{A}}_{\max}}\right|\log^{2}(2\left|{{\mathcal{S}}}\right|/\delta)}{N}},$
where $C_{\max}=\max_{\mu,\nu}\max\left\{\max_{\nu^{\star}\in\mathrm{br}(\mu)}\left\|{\frac{d^{\mu^{\operatorname{E}},\nu^{\star}}}{\rho}}\right\|_{\infty},\max_{\mu^{\star}\in\mathrm{br}(\nu)}\left\|{\frac{d^{\mu^{\star},\nu^{\operatorname{E}}}}{\rho}}\right\|_{\infty}\right\}$ and we set $\rho=d^{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}.$
@@
[…]
@@
We note that $\max_{\mu\in\mathrm{br}(\widehat{\nu})}$ contains the random variable $\hat{\nu}$ but can be upper bounded with the best response to with respect to the general policy class $\max_{\nu\in\Pi}\max_{\mu\in\mathrm{br}(\nu)}$. Last, we do a change of measure to get the expectation with respect to the expert policy pair. Then, we bound the ratio of the state visitation distribution and bound the expectation with concentration inequalities to obtain with probability of at least $1-\delta$
$V^{\mu^{\star},\widehat{\nu}}(s_{0})-V^{\mu^{\star},\nu^{\operatorname{E}}}(s_{0})\leq\frac{8}{(1-\gamma)^{2}}\max_{\nu\in\Pi}\max_{\mu\in\mathrm{br}(\nu)}\left\|\frac{d^{\mu,\nu^{\operatorname{E}}}}{d^{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}}\right\|_{\infty}\sqrt{\frac{\left|{{\mathcal{S}}}\right|\left|{{\mathcal{B}}}\right|\log^{2}(2\left|{{\mathcal{S}}}\right|/\delta)}{N}}.$
Analogous calculations for the second player complete the proof. The full proof is given in Appendix˜C.∎
We now discuss several important implications of the derived theorem, particularly focusing on the quantity ${\mathcal{C}}_{\max}$, referred to as the all policy deviation concentrability coefficient (see, e.g., (Cui and Du, 2022a; Zhong et al., 2022)). Intuitively, the theorem indicates that if a best response against a possible recovered policy shifts the support of the state visitation distribution away from the one induced by the observed Nash equilibrium, the corresponding objective becomes unbounded.
@@
While restricting, this requirement is weaker than a uniform lower bound on the equilibrium state occupancy measure assumed by Tang et al. (2024), that is $d^{\mu_{E},\nu_{E}}\geq\beta$. In particular, it always holds that $\mathcal{C}_{\max}\leq\beta^{-1}$.
Next, we introduce a lower bound on ${\mathcal{C}}_{\max}$, that intuitively describes how different the coverage of different Nash equilibrium policies can be. Formally, we introduce ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}}):=\max\left\{\max_{\mu\in\mathrm{br}(\nu^{\operatorname{E}})}\left\|\frac{d^{\mu,\nu^{\operatorname{E}}}}{d^{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}}\right\|_{\infty},\max_{\nu\in\mathrm{br}(\mu^{\operatorname{E}})}\left\|\frac{d^{\mu^{\operatorname{E}},\nu}}{d^{\mu^{\operatorname{E}},\nu^{\operatorname{E}}}}\right\|_{\infty}\right\},$ which we refer to as the expert policy deviation concentrability coefficient. Clearly, we have that ${\mathcal{C}}_{\max}\geq{\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$ as the latter only consider the best responses to the Nash equilibrium policies.
Unfortunately, we show that even for this smaller quantity there exists a zero-sum Markov game in which $\mathcal{C}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$ is unbounded, and we show that in such a game, no non-interactive algorithm can recover a Nash profile even with an infinite amount of data. We present this result in the next section. Note, that it remains an open question if it is possible to have a bounded ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$ but an unbounded ${\mathcal{C}}_{\max}.$ We discuss this further in Section˜6.
@@
3.1 Necessity of $\mathcal{C}(\mu_{E},\nu_{E})$ in non-interactive MAIL
In this section, we provide the negative result, that a Markov Game exists, such that the expert deviation concentrability coefficient of Theorem˜3.1 is unbounded.
The first hardness results to minimize the Nash Gap in Multi-Agent Imitation Learning were derived by  Tang et al. (2024, Thm. 4.3). Next, we will give a stronger result, showing that even in the case of full knowledge of the transition model and perfect recovery of the state visitation distribution of the expert, the Nash gap is of the order $(1-\gamma)^{-1}$. A detailed discussion on the difference between the following result and the one obtained by Tang et al. (2024, Thm. 4.3) can be found in Appendix˜I. An illustration of the Zero-Sum Markov game can be found in Fig.˜1 and the full proof in Appendix˜D.